# Undefinability of Approximation of 2-to-2 Games

### v1

@@ 5.3 Graph Colouring
It is shown in [16, Lemma 10] that such matrices exist. Having fixed such a $T$, we can define the reduction. The reduction takes an instance $(V,E,\Sigma,\Phi)$, a $2\leftrightarrow 2$-game with $\Sigma=\{0,\ldots,2q-1\}$, to a graph $(V^{\prime},E^{\prime})$. The nodes in $V^{\prime}$ are tuples $(v,x_{0},\dots,x_{2q-1})$, where $v\in V$ and each $x_{i}\in\{0,1,2,3\}$. There is an edge in $E^{\prime}$ between $(u,x_{0},\dots,x_{2q-1})$ and $(v,y_{0},\dots,y_{2q-1})$ if there is some constraint $e\in E$ linking $u$ and $v$, such that $\Phi(e)=D_{\pi_{1},\pi_{2}}$ as defined in Section 3.5) such that:
@@
This condition, combined with the second condition on $T$, ensures that given a labeling $\chi:V\rightarrow\Sigma$ that satisfies all constraints, the colouring of $V^{\prime}$ that assigns $x_{\chi(v)}$ to the node $(v,x_{0},\dots,x_{2q-1})$ is a proper $4$-colouring of $(V^{\prime},E^{\prime})$. On the other hand, if $(V,E,\Sigma,\Phi)$ is not $\delta$-satisfiable, $(V^{\prime},E^{\prime})$ does not have an independent set of relative size $\delta^{\prime}$, for a value of $\delta^{\prime}$ that tends to 00 as $\delta$ does. Hence the graph is not $\frac{1}{\delta^{\prime}}$-colourable. The proof of this can be found in [16, Section 3.3].
@@
This means that the graph transformation $(\mathrm{sym}\circ\mathrm{arc}\circ\mathrm{arc}\circ\mathrm{dir})$ is a reduction from $\mathrm{GapColour}(4,2^{2^{t}})$ to $\mathrm{GapColour}(3,t)$.
@@
$\phi_{E}(x,y):=\exists zC(z)\land\Phi_{\Pi}(x_{3},y_{3},z)\land\Pi(x,y).$
@@ 6 Conclusion
We have shown that the reductions involved in the proof of the celebrated proof by Khot, Minzer and Safra of the $2$-to-$2$ games theorem can all be implemented as interpretations in first-order logic. This means that the $\mathrm{NP}$-hardness they establish of separating nearly satisfiable instances from highly unsatisfiable ones can be turned into an unconditional inseparability result in $\mathrm{FPC}$. Moreover, the result is achieved with perfect completeness: it is impossible to separate with an $\mathrm{FPC}$ sentence the fully satisfiable $2$-to-$2$ games from those that are highly unsatisfiable.
From this result we are able to derive a number of consequences, the most striking of which is that it is impossible to separate with an $\mathrm{FPC}$ sentence the graphs that are $3$-colourable from those that are not $t$-colourable for any constant $t$. The $\mathrm{NP}$-hardness of such a separation is only conjectured for values $t$ larger than $5$. Moreover, oure result strengthens the undefinabiliety resutls of Atserias and Dalmau [5]. We also obtain strong $\mathrm{FPC}$ undefinability results for approximation of unique games. In terms of approximation ratios these are an improvement over those of Tucker-Foltz [28]. However, the latter results were obtained for simple games while ours are for games with multiple edges.

### v2

@@ 5.3 Graph Colouring
It is shown in [18, Lemma 10] that such matrices exist. Having fixed such a $T$, we can define the reduction. The reduction takes an instance $(V,E,\Sigma,\Phi)$, a $2\leftrightarrow 2$-game with $\Sigma=\{0,\ldots,2q-1\}$, to a graph $(V^{\prime},E^{\prime})$. The nodes in $V^{\prime}$ are tuples $(v,x_{0},\dots,x_{2q-1})$, where $v\in V$ and each $x_{i}\in\{0,1,2,3\}$. There is an edge in $E^{\prime}$ between $(u,x_{0},\dots,x_{2q-1})$ and $(v,y_{0},\dots,y_{2q-1})$ if there is some constraint $e\in E$ linking $u$ and $v$, such that $\Phi(e)=D_{\pi_{1},\pi_{2}}$ as defined in Section 3.5) such that:
@@
This condition, combined with the second condition on $T$, ensures that given a labeling $\chi:V\rightarrow\Sigma$ that satisfies all constraints, the colouring of $V^{\prime}$ that assigns $x_{\chi(v)}$ to the node $(v,x_{0},\dots,x_{2q-1})$ is a proper $4$-colouring of $(V^{\prime},E^{\prime})$. On the other hand, if $(V,E,\Sigma,\Phi)$ is not $\delta$-satisfiable, $(V^{\prime},E^{\prime})$ does not have an independent set of relative size $\delta^{\prime}$, for a value of $\delta^{\prime}$ that tends to 0 as $\delta$ does. Hence the graph is not $\frac{1}{\delta^{\prime}}$-colourable. The proof of this can be found in [18, Section 3.3].
@@
This means that the graph transformation $(\mathrm{sym}\circ\mathrm{arc}\circ\mathrm{arc}\circ\mathrm{dir})$ is a reduction from
$\mathrm{GapColour}(4,2^{2^{t}})$ to $\mathrm{GapColour}(3,t)$.
@@
$\phi_{E}(x,y):=\exists zC(z)\land\bigvee_{\Pi}[\Phi_{\Pi}(x_{3},y_{3},z)\land\Pi(x,y)].$
@@ 6 Conclusion
We have shown that the reductions involved in the proof of the celebrated proof by Khot, Minzer and Safra of the 22-to-22 games theorem can all be implemented as interpretations in first-order logic. This means that the $\mathrm{NP}$-hardness they establish of separating nearly satisfiable instances from highly unsatisfiable ones can be turned into an unconditional inseparability result in $\mathrm{FPC}$. Moreover, the result is achieved with perfect completeness: it is impossible to separate with an $\mathrm{FPC}$ sentence the fully satisfiable 22-to-22 games from those that are highly unsatisfiable.
From this result we are able to derive a number of consequences, the most striking of which is that it is impossible to separate with an $\mathrm{FPC}$ sentence the graphs that are $3$-colourable from those that are not $t$-colourable for any constant $t$. The $\mathrm{NP}$-hardness of such a separation is only conjectured for values $t$ larger than $5$. Moreover, our result strengthens the undefinabiliety results of Atserias and Dalmau [6]. We also obtain strong $\mathrm{FPC}$ undefinability results for approximation of unique games. In terms of approximation ratios these are an improvement over those of Tucker-Foltz [29]. However, the latter results were obtained for simple games while ours are for games with multiple edges.